# Flavor Patterns of Fundamental Particles from Quantum Entanglement?

### v1

@@ I Introduction
The Standard Model (SM) of particle physics—which provides our most precise understanding of microscopic physics in our Universe—has 19 input parameters related to the masses, mixings, and couplings between quarks, leptons, gauge bosons, and the Higgs boson. A grand challenge in fundamental physics is to understand the origins of, and relations between, these 19 parameters. The most notable example of possible parameter reduction appears in the gauge sector, where the strong, electroweak, and hypercharge gauge couplings could have a common origin in grand unified theories [1, 2, 3]. Just as the dizzying spectrum of strongly-interacting hadrons were explained in terms of constituent quarks and gluons in quantum chromodynamics (QCD) [4, 5, 6, 7, 8, 9], the puzzle of SM parameters could be resolved by finding new high-energy structures.
@@
Entanglement is a core phenomenon in quantum mechanics, where measurement outcomes are correlated beyond classical expectations. In particle physics, entanglement is so ubiquitous that we often take it for granted, but every neutral pion decay to two photons ($\pi_{0}\to\gamma\gamma$) is effectively a mini Einstein–Podolsky–Rosen experiment. In the context of SM scattering processes, though, the study and quantification of entanglement in its own right has only begun relatively recently [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In terms of predicting particle properties from entanglement, the first paper we are aware of is Ref. [27] which showed that maximizing helicity entanglement yields a reasonable prediction for the Weinberg angle $\theta_{W}$, which controls the mixing between electroweak bosons.
Our study of entanglement for flavor physics was motivated by recent work on the interplay between entanglement in particle scattering and the emergence of global symmetries at low energies. Refs. [19, 21] showed that the suppression of spin-entanglement in non-relativistic $np$ scattering leads to Wigner’s $SU(4)_{sm}$ spin-flavor symmetry [28] and to the Schrödinger symmetry [29, 30]. Ref. [22] showed that the suppression of flavor-entanglement in scattering of octet baryons leads to enlarged symmetries consistent with phenomenology. Ref. [25] examined the entanglement produced by scattering in a generic two-Higgs doublet model, finding that vanishing entanglement leads to the maximal symmetry $SO(8)$. [^1] Ref. [31] recently highlighted the importance of analyzing the entanglement properties of all different scattering channels between the external particles. Based on these previous studies, we expected that suppressing entanglement for scattering involving flavor would again lead to enhanced symmetries, for example an enhanced $SU(3)$ flavor symmetry rotating all three quark generations.
@@
Regardless of whether entanglement minimization is a principle of nature, the appearance of any prediction for a hierarchical CKM matrix that simultaneously predicts an anarchic PMNS matrix is cause for celebration (or alarm, depending on your perspective). The quark and lepton flavor sectors of the SM have been notoriously difficult to decipher. In pursuit of an underlying organizing principle, spurion analysis has been employed, often motivated by numerical relations among the fermion masses and CKM elements [37, 38, 39, 40]. Traditionally, these investigations have led to the proposal of flavor symmetries at high energies that appear broken in the low-energy limit [41, 42, 32, 33, 43, 44]. For example, in Froggatt–Nielson (FN) models [41, 45, 46], the structure of the CKM matrix is determined by $U(1)$ symmetries under which different fermion flavors have different charges. Such models can help explain why the Cabibbo angle $\theta_{C}$ is small but non-zero, since it arises from a ratio of mass scales. This kind of spurion analysis, however, is not well matched to the structure of the PMNS matrix. To our knowledge, this is the first time the differing CKM and PMNS structures have arisen from a common mechanism that does not invoke symmetries or symmetry breaking.
@@ II.1 Flavor Hilbert Space
These correspond respectively to the quark flavors up, charm, top, down, strange, and bottom. The anti-particles belong to the complex conjugate Hilbert spaces $H_{\bar{u}}$ and $H_{\bar{d}}$, which are however isomorphic to $H_{u}$ and $H_{d}$. Similarly, we can define Hilbert spaces for the leptons $H_{\ell}$ and neutrinos $H_{\nu}$.
@@ II.2 Entangling Power
Next, we have to choose the states to average over in Eq. (11). In the absence of fermion masses, the CKM and PMNS matrices can be diagonalized by doing a chiral rotation. This in turn means that Eq. (11) will have no sensitivity to flavor mixing angles in the massless limit, unless the average depends on a preferred choice of flavor frame. We therefore choose to average over product states of definite fermion generation:
@@ II.3 Scattering Channels and Kinematics
where the subscripts corresponds to the initial state of the scattering channel.
@@ III Entangling Power of Electroweak Interactions
We start with the case of two quark flavors to gain intuition, and then present three-generation results. The relevant scattering amplitudes are reviewed in App. C.
@@ III.1 The Cabibbo Angle
To convert this amplitude to the perpendicular entangling power, we fix the kinematics to $t=u=-\frac{s}{2}$ and normalize -the final state wav in each |$\ket{ij}_{ud}$ channel separately. In the $4$-dimensional $H_{f}$ space for two generations, following the definition in Eq. (49):
@@
Further inspecting Eq. (III.1), we see that the Weinberg angle $\theta_{W}$ is the only other SM coupling that affects the value of the Cabibbo angle $\theta_{C}^{\rm min}$ at the entangling power minimum. Roughly speaking, the $W$-induced charged-current process induces flavor entanglement while the $\gamma/Z$-induced neutral-current process suppresses it. When taking both processes into account, it is this relative imbalance that yields a non-trivial flavor-mixing angle.
Going into more detail, we study the dependence of $\theta_{C}^{\rm min}$ on $\theta_{W}$ in Fig. 2b. As $\sin\theta_{W}\to 0$, the contribution from flavor-changing $W$-boson exchange becomes larger, therefore minimizing entanglement requires a smaller (though non-vanishing) value for $\theta_{C}^{\rm min}$. The reason why $\theta_{C}^{\rm min}$ is still non-zero at $\sin\theta_{W}=0$ arises from the initial-state flavor averaging procedure. From Eq. (24), we see that the CKM factor takes different values for flavor-preserving versus flavor-changing scattering:
@@
The minimum is then a compromise between the flavor-changing cases that generate entanglement and the flavor-preserving case that does not but still destructively interferes with $Z$-boson induced channels (see the relative minus sign in Eq. (22)). In the opposite limit $\sin\theta_{W}\to 1$ of vanishing weak isospin, the $Z$-boson exchange dominates, so the value of $\theta_{C}$ is determined primarily by the suppression of the destructive interference in the flavor-preserving channel.
@@ III.2 Towards the Full CKM Matrix
Figure 3: Electroweak entanglement in the quark sector as a function of the scattering center-of-mass energy $\sqrt{s}$. (a) The perpendicular entangling power after minimizing over the CKM angles, considering the two scattering channels $ud\to ud$ and $u\bar{d}\to u\bar{d}$ separately. (b) The CKM angles at the entangling power minimum, with solid lines corresponding to the channel that minimizes $\mathcal{E}_{\rm min}^{\perp}$ at each $\sqrt{s}$ and dashed lines to the other channel for reference. The CP-violating phase $\delta_{\rm CKM}$ has no impact on the minimization.
@@
Since we do not have a principle to set the scattering energy, it is interesting to study the entangling behavior as a function of $\sqrt{s}$. At energies far below the electroweak scale, the flavor-preserving photon exchange dominates, so entanglement suppression no longer enforces a flavor-diagonal coupling of the charged current, leading to maximal mixing between the first two generations. Above but close to the top quark mass, $\sqrt{s}\gtrsim m_{t}$, the mass differences between the fermions become negligible and we get roughly equal mixing angles around $4^{\circ}$. In the high energy limit with $\sqrt{s}\gg m_{t}$, the $u\bar{d}\to u\bar{d}$ channel becomes the one that minimizes the entangling power, which prefers no quark mixing.
@@ III.3 Towards the Full PMNS Matrix
Figure 4: Electroweak entanglement in the lepton sector as a function of the scattering center-of-mass energy $\sqrt{s}$. (a) The perpendicular entangling power after minimizing over the PMNS angles, considering the two scattering channels $\nu\ell\to\nu\ell$ and $\nu\bar{\ell}\to\nu\bar{\ell}$ separately. The inset focuses on the region around the minimizing energy to illustrate the effect of the CP-violating phase for the benchmarks $\delta_{\rm{PMNS}}^{\rm{NO}}=197^{\circ}$ (solid line) and $\delta_{\rm{PMNS}}^{\rm{IO}}=286^{\circ}$ (dotted line). (b) The PMNS angles at the entangling power minimum, with solid lines corresponding to the $\nu\ell\to\nu\ell$ channel that minimizes $\mathcal{E}_{\rm min}^{\perp}$ after fixing the CP-violating phase at $\delta_{\rm{PMNS}}^{\rm{NO}}=197^{\circ}$, dotted lines to the same but at $\delta_{\rm{PMNS}}^{\rm{IO}}=286^{\circ}$, and dashed lines to minimizing the $\nu\bar{\ell}\to\nu\bar{\ell}$ channel for reference.
@@
There is an intriguing interplay between $\delta_{\rm{PMNS}}$ and the neutrino mass hierarchy. Global analyses [51] provide two different best-fit PMNS points depending normal ordering (NO, $m_{\nu,1}<m_{\nu,2}<m_{\nu,3}$) versus inverted ordering (IO, $m_{\nu,3}<m_{\nu,1}<m_{\nu,2}$) of the neutrino masses. The NO best fit of $\delta_{\rm{PMNS}}^{\rm{exp,NO}}={197^{\circ}}^{+41^{\circ}}_{-25^{\circ}}$ is consistent with entangling minimum at $\pi$, while the IO best fit of $\delta_{\rm{PMNS}}^{\rm{exp,IO}}={286^{\circ}}^{+27^{\circ}}_{-32^{\circ}}$ is close to the maximum at $23\pi/$ 2. As shown with the dotted lines in Fig. 4a, the difference in the value of $\mathcal{E}_{\nu\ell}^{\perp}$ between the minimum (solid black) and maximum (dotted black) is relatively small and could be modified by higher-order effects. Taken seriously, though, this analysis would further “predict” normal ordering of neutrinos with minimal CP violation.
The values of the PMNS angles that minimize the entangling power are shown in Fig. 4b. For most of the range $m_{\tau}\lesssim\sqrt{s}\lesssim m_{Z}$, the minimum happens when all of the angles are sizeable, in agreement with the qualitative experimental picture of lepton mixing. The energy scale that minimizes $\mathcal{E}_{\nu\ell}^{\perp}$ in this range is $\sqrt{s}\approx 30~{}\rm GeV$. Fixing this scale and and setting $\delta_{\rm PMNS}$ to the NO best fit value, we find the following PMNS angles:
@@
Compared to the quark case, it is somewhat easier to understand the qualitative behavior as a function of $\sqrt{s}$. Because the photon channel does not contribute, the only competition is between diagrams with $W$ or $Z$ exchange, which is present at all energy scales. Balancing these diagrams happens when mixing angles are of comparable same size. Below the tau threshold $\sqrt{s}\lesssim m_{\tau}$, the three neutrinos scatter against two charged leptons, and balancing flavor-preserving against flavor-violating scattering requires two nearly maximal angles. Far enough above the tau threshold, $\sqrt{s}\gg m_{\tau}$, all of the mixing angles need to be around the same size to minimize the entangling power.
@@ IV Conclusions
Ultimately, we want to know: is this all just a numerical coincidence, or could minimization of quantum entanglement really be a fundamental principle of nature? To answer this question, we encourage further explorations of the consequences of entanglement in various scattering processes, both in and beyond the SM. While previous studies have focused on the emergence of symmetries, our analysis shows that entanglement suppression can lead to predictions for symmetry-breaking effects, suggesting that other kinds of consequences could arise from this simple principle. Though speculative, this line of research highlights the potential of injecting concepts from quantum information theory into high energy physics.
@@ Acknowledgements.
We thank Darius Faroughy, Yuval Grossman, Gino Isidori, Ian Low, Grant Remmen, Olcyr Sumensari, Raymond Volkas, and Carlos Wagner for useful discussions and appropriately skeptical feedback on our work. This work was supported by the Office of High Energy Physics of the U.S. Department of Energy (DOE) under Grant No. DE-SC0012567, and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab (FNAL 20-17). J.T. is additionally supported by the DOE Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under Contract No. DE-SC0012704, by the Simons Foundation through Investigator grant 929241, and his work was performed in part at the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452. S.T. was additionally supported by the Swiss National Science Foundation - project n. P500PT_203156.

### v2

@@ I Introduction
The Standard Model (SM) of particle physics—which provides our most precise understanding of microscopic physics in our Universe—has 19 input parameters related to the masses, mixings, and couplings between quarks, leptons, gauge bosons, and the Higgs boson. A grand challenge in fundamental physics is to understand the origins of, and relations between, these 19 parameters. The most notable example of possible parameter reduction appears in the gauge sector, where the strong, electroweak, and hypercharge gauge couplings could have a common origin in grand unified theories [1, 2, 3]. Just as the dizzying spectrum of strongly-interacting hadrons was explained in terms of constituent quarks and gluons in quantum chromodynamics (QCD) [4, 5, 6, 7, 8, 9], the puzzle of SM parameters could be resolved by finding new high-energy structures.
@@
Entanglement is a core phenomenon in quantum mechanics, where measurement outcomes are correlated beyond classical expectations. In particle physics, entanglement is so ubiquitous that we often take it for granted, but every neutral pion decay to two photons ($\pi_{0}\to\gamma\gamma$) is effectively a mini Einstein–Podolsky–Rosen experiment. In the context of SM scattering processes, though, the study and quantification of entanglement in its own right has only begun relatively recently [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. In terms of predicting particle properties from entanglement, the two papers we are aware of are: i) Ref. [29] which showed that maximizing helicity entanglement yields a reasonable prediction for the Weinberg angle $\theta_{W}$ that controls the mixing between electroweak bosons, and ii) Ref. [30] which showed that minimizing entanglement in neutrino oscillations yields a prediction for the CP-violating phase of the PMNS matrix.
Our study of entanglement for flavor physics was motivated by recent work on the interplay between entanglement in particle scattering and the emergence of global symmetries at low energies. Refs. [19, 21, 22] showed that the suppression of spin-entanglement in non-relativistic baryon scattering leads to enlarged symmetries, such as Wigner’s $SU(4)_{sm}$ spin-flavor symmetry [31] for two flavors, $SU(16)$ for three, and the Schrödinger conformal invariance [32, 33]. Ref. [25] examined the entanglement produced by scattering in a generic two-Higgs doublet model, finding that vanishing entanglement leads to the maximal symmetry $SO(8)$. [^1] Ref. [34, 35] recently highlighted the importance of analyzing the entanglement properties of all different scattering channels between the external particles. Based on these previous studies, we expected that suppressing entanglement for scattering involving flavor would again lead to enhanced symmetries, for example an enhanced $SU(3)$ flavor symmetry rotating all three quark generations.
@@
Regardless of whether entanglement minimization is a principle of nature, the appearance of any prediction for a hierarchical CKM matrix that simultaneously predicts an anarchic PMNS matrix is cause for celebration (or alarm, depending on your perspective). The quark and lepton flavor sectors of the SM have been notoriously difficult to decipher. In pursuit of an underlying organizing principle, spurion analysis has been employed, often motivated by numerical relations among the fermion masses and CKM elements [41, 42, 43, 44]. Traditionally, these investigations have led to the proposal of flavor symmetries at high energies that appear broken in the low-energy limit [45, 46, 36, 37, 47, 48, 49]. For example, in Froggatt–Nielsen (FN) models [45, 50, 51], the structure of the CKM matrix is determined by $U(1)$ symmetries under which different fermion flavors have different charges. Such models can help explain why the Cabibbo angle $\theta_{C}$ is small but non-zero, since it arises from a ratio of mass scales. This kind of spurion analysis, however, is not well matched to the structure of the PMNS matrix. To our knowledge, this is the first time the differing CKM and PMNS structures have arisen from a common mechanism that does not invoke symmetries or symmetry breaking.
@@ II.1 Flavor Hilbert Space
These correspond, respectively, to the quark flavors up, charm, top, down, strange, and bottom. The anti-particles belong to the complex conjugate Hilbert spaces $H_{\bar{u}}$ and $H_{\bar{d}}$, which are, however, isomorphic to $H_{u}$ and $H_{d}$. Similarly, we can define Hilbert spaces for the leptons $H_{\ell}$ and neutrinos $H_{\nu}$.
@@ II.2 Entangling Power
Next, we have to choose the states to average over in Eq. (11). In the absence of fermion masses, the CKM and PMNS matrices can be diagonalized by doing a chiral rotation. This, in turn, means that Eq. (11) will have no sensitivity to flavor mixing angles in the massless limit, unless the average depends on a preferred choice of flavor frame. We therefore choose to average over product states of definite fermion generation:
@@ II.3 Scattering Channels and Kinematics
where the subscripts correspond to the initial state of the scattering channel.
@@ III Entangling Power of Electroweak Interactions
We start with the case of two quark generations to gain intuition, and then present three-generation results. The relevant scattering amplitudes are reviewed in App. C.
@@ III.1 The Cabibbo Angle
To convert this amplitude to the perpendicular entangling power, we fix the kinematics to $t=u=-\frac{s}{2}$ and normalize the final-state wavefunctions in each |$\ket{ij}_{ud}$ channel separately. In the $4$-dimensional $H_{f}$ space for two generations, following the definition in Eq. (49):
@@
We can understand how the non-trivial flavor-mixing angle arises from the minimization by observing the following relative imbalance: the $W$-induced charged-current process generates both flavor-changing and flavor-conserving effects, while the $\gamma/Z$-induced neutral-current process only flavor-conserving effects. In fact, according to Eq. (24) the CKM factor that appears in the charged-current term takes different values depending on whether the flavor of each quark has changed or not:
@@
When taking both processes into account, in case the initial quarks are of the same generation the flavor-changing effects enhance flavor entanglement, while in case the initial quarks are of different generations they suppress it (see Eqs. (30) and (31)).
Further inspecting Eq. (III.1), we see that the Weinberg angle $\theta_{W}$ is the only other SM coupling that affects the value of the Cabibbo angle $\theta_{C}^{\rm min}$ at the entangling power minimum. We study this dependence of $\theta_{C}^{\rm min}$ on $\theta_{W}$ in Fig. 2b. As $\sin\theta_{W}\to 0$, the contribution from the flavor-changing $W$-boson exchange becomes larger, and thus minimizing entanglement requires a smaller (though non-vanishing) value for $\theta_{C}^{\rm min}$. The minimum is then a compromise between the flavor-changing effects of the $W$ boson that generate entanglement and those that suppress it, as explained above. In the opposite limit $\sin\theta_{W}\to 1$, the $Z$-boson exchange dominates and Eq. (III.1) simplifies to $\mathcal{E}_{ud}^{\perp}=(\cos^{4}\theta_{C}+\sin^{4}\theta_{C})/(2y^{2})$. Therefore $\theta_{C}=\pi/4$ becomes the preferred minimum.
@@ III.2 Towards the Full CKM Matrix
Figure 3: Electroweak entanglement in the quark sector as a function of the scattering center-of-mass energy $\sqrt{s}$. (a) The perpendicular entangling power after minimizing over the CKM angles, considering the two scattering channels $ud\to ud$ and $u\bar{d}\to u\bar{d}$ separately. (b) The CKM angles at the entangling power minimum, with solid lines corresponding to the channel that minimizes $\mathcal{E}^{\perp}$ at each $\sqrt{s}$ and dashed lines to the other channel for reference. The CP-violating phase $\delta_{\rm CKM}$ has no impact on the minimization.
@@
Since we do not have a principle to set the scattering energy, it is interesting to study the entangling behavior as a function of $\sqrt{s}$. At energies far below the electroweak scale, the flavor-preserving photon exchange dominates, so entanglement suppression no longer enforces a flavor-diagonal coupling of the charged current, leading to maximal mixing between the first two generations. Above but close to the top quark mass, $\sqrt{s}\gtrsim m_{t}$, the mass differences between the fermions become negligible and we get roughly equal mixing angles around $4^{\circ}$. In the high-energy limit with $\sqrt{s}\gg m_{t}$, the $u\bar{d}\to u\bar{d}$ channel becomes the one that minimizes the entangling power, which prefers no quark mixing.
@@ III.3 Towards the Full PMNS Matrix
Figure 4: Electroweak entanglement in the lepton sector as a function of the scattering center-of-mass energy $\sqrt{s}$. (a) The perpendicular entangling power after minimizing over the PMNS angles, considering the two scattering channels $\nu\ell\to\nu\ell$ and $\nu\bar{\ell}\to\nu\bar{\ell}$ separately. The inset focuses on the region around the minimizing energy to illustrate the effect of the CP-violating phase for the benchmarks $\delta_{\rm{PMNS}}^{\rm{NO}}=197^{\circ}$ (solid line) and $\delta_{\rm{PMNS}}^{\rm{IO}}=286^{\circ}$ (dotted line). (b) The PMNS angles at the entangling power minimum, with solid lines corresponding to the $\nu\ell\to\nu\ell$ channel that minimizes $\mathcal{E}^{\perp}$ after fixing the CP-violating phase at $\delta_{\rm{PMNS}}^{\rm{NO}}=197^{\circ}$, dotted lines to the same but at $\delta_{\rm{PMNS}}^{\rm{IO}}=286^{\circ}$, and dashed lines to minimizing the $\nu\bar{\ell}\to\nu\bar{\ell}$ channel for reference.
@@
There is an intriguing interplay between $\delta_{\rm{PMNS}}$ and the neutrino mass hierarchy. Global analyses [56] provide two different best-fit PMNS points depending on normal ordering (NO, $m_{\nu,1}<m_{\nu,2}<m_{\nu,3}$) versus inverted ordering (IO, $m_{\nu,3}<m_{\nu,1}<m_{\nu,2}$) of the neutrino masses. The NO best fit of $\delta_{\rm{PMNS}}^{\rm{exp,NO}}={197^{\circ}}^{+41^{\circ}}_{-25^{\circ}}$ is consistent with the entangling minimum at $\pi$, while the IO best fit of $\delta_{\rm{PMNS}}^{\rm{exp,IO}}={286^{\circ}}^{+27^{\circ}}_{-32^{\circ}}$ is close to the maximum at $23\pi/$ 2. As shown with the dotted lines in Fig. 4a, the difference in the value of $\mathcal{E}_{\nu\ell}^{\perp}$ between the minimum (solid black) and maximum (dotted black) is relatively small and could be modified by higher-order effects. Taken seriously, though, this analysis would further “predict” normal ordering of neutrinos with minimal CP violation. [^10] Note that Ref. [30] also comes to the same conclusion from an entanglement minimization principle applied to neutrino oscillations. Given the different origin of entanglement generation from the one considered in this work, we do not discern at present any underlying connection between these two observations.
The values of the PMNS angles that minimize the entangling power are shown in Fig. 4b. For most of the range $m_{\tau}\lesssim\sqrt{s}\lesssim m_{Z}$, the minimum happens when all of the angles are sizeable, in agreement with the qualitative experimental picture of lepton mixing. The energy scale that minimizes $\mathcal{E}_{\nu\ell}^{\perp}$ in this range is $\sqrt{s}\approx 30~{}\rm GeV$. Fixing this scale and setting $\delta_{\rm PMNS}$ to the NO best fit value, we find the following PMNS angles:
@@
Compared to the quark case, it is somewhat easier to understand the qualitative behavior as a function of $\sqrt{s}$. Because the photon channel does not contribute, the only competition is between diagrams with $W$ or $Z$ exchange, which is present at all energy scales. Balancing these diagrams happens when mixing angles are of comparable size. Below the tau threshold $\sqrt{s}\lesssim m_{\tau}$, the three neutrinos scatter against two charged leptons, and balancing flavor-preserving against flavor-violating scattering requires two nearly maximal angles. Far enough above the tau threshold, $\sqrt{s}\gg m_{\tau}$, all of the mixing angles need to be around the same size to minimize the entangling power.
@@ IV Conclusions
Ultimately, we want to know: is this all just a numerical coincidence, or could minimization of quantum entanglement really be a fundamental principle of nature? To answer this question, we encourage further explorations of the consequences of entanglement in various scattering processes, both in and beyond the SM. While previous studies have focused on the emergence of symmetries, our analysis shows that entanglement suppression can lead to predictions for symmetry-breaking effects, suggesting that other kinds of consequences could arise from this simple principle. Though speculative, this line of research highlights the potential of injecting concepts from quantum information theory into high-energy physics.
@@ Acknowledgements.
We thank Gilly Elor, Darius Faroughy, Yuval Grossman, Gino Isidori, Ian Low, Grant Remmen, Olcyr Sumensari, Raymond Volkas, and Carlos Wagner for useful discussions and appropriately skeptical feedback on our work. This work was supported by the Office of High Energy Physics of the U.S. Department of Energy (DOE) under Grant No. DE-SC0012567, and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab (FNAL 20-17). J.T. is additionally supported by the DOE Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under Contract No. DE-SC0012704, by the Simons Foundation through Investigator grant 929241, and his work was performed in part at the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452. S.T. was additionally supported by the Swiss National Science Foundation - project n. P500PT_203156.